# PAR-AdvGAN: Improving Adversarial Attack Capability with Progressive Auto-Regression AdvGAN

### v1

@@ Abstract
Deep neural networks have demonstrated remarkable performance across various domains. However, they are vulnerable to adversarial examples, which can lead to erroneous predictions. Generative Adversarial Networks (GANs) can leverage the generators and discriminators model to quickly produce high-quality adversarial examples. Since both modules train in a competitive and simultaneous manner, GAN-based algorithms like AdvGAN can generate adversarial examples with better transferability compared to traditional methods. However, the generation of perturbations is usually limited to a single iteration, preventing these examples from fully exploiting the potential of the methods. To tackle this issue, we introduce a novel approach named Progressive Auto-Regression AdvGAN (PAR-AdvGAN). It incorporates an auto-regressive iteration mechanism within a progressive generation network to craft adversarial examples with enhanced attack capability. We thoroughly evaluate our PAR-AdvGAN method with a large-scale experiment, demonstrating its superior performance over various state-of-the-art black-box adversarial attacks, as well as the original AdvGAN. Moreover, PAR-AdvGAN significantly accelerates the adversarial example generation, i.e., achieving the speeds of up to 335.5 frames per second on Inception-v3 model, outperforming the gradient-based transferable attack algorithms. Our code is available at: https://anonymous.4open.science/r/PAR-01BF/
@@ 1 Introduction
Deep neural networks (DNNs) are widely used in different real-world applications, i.e., image classification (Li et al., 2020), emotional analysis (Qiang et al., 2020), and item recommendations (Pan et al., 2020). DNNs demonstrate human-surpassed performance when properly trained. However, DNNs can be vulnerable to adversarial examples crafted by attackers (Szegedy et al., 2013; Ma et al., 2021; Deng et al., 2020), which is a concern in safety-critical scenarios. Thus, a practical approach is to develop effective attack algorithms that can assess the robustness of DNNs against adversarial attacks at an early stage, ultimately enhancing model safety.
@@
Currently, both white-box and black-box attack algorithms, such as gradient-based methods like FGSM (Goodfellow et al., 2014b), NAA (Zhang et al., 2022), SSA (Long et al., 2022), and optimization-based approaches such as PGD (Madry et al., 2017) and C&W (Carlini & Wagner, 2017), require continuous computation of the model’s gradient information throughout the attack process. However, they all require extensive running time. Generative Adversarial Networks (GANs) (Goodfellow et al., 2014a) have demonstrated promising results for realistic sample generation by leveraging both generator and discriminator for training (Karras et al., 2021; Chang et al., 2022; Haidar & Rezagholizadeh, 2019; Croce et al., 2020). While the generator constructs high-quality examples, a discriminator learn to distinguish the original and generated examples. Furthermore, once the generator is trained, there will be no additional gradient computation for input examples.
@@
As an early GAN-based model, AdvGAN incorporates a perturbation, denoted as $G(x)$, into the original image instance $x$ for attack (Xiao et al., 2018). AdvGAN aims to obtain the manipulated image $x+G(x)$ from the original instance $x$ through the discriminator. To achieve high attack success rates in both white-box and black-box attacks, AdvGAN introduces an adversarial loss on top of GANs loss, ensuring the adversarial image is generated in a direction more effective for adversarial attacks. Additionally, it employs hinge loss to limit perturbation range, thereby preventing significant deviations between the adversarial and original images. Subsequently, AdvGAN++ (Jandial et al., 2019) further enhances the attack success rate by utilizing latent features instead of input image instances $x$. It optimizes the latent features during adversarial examples generation.
@@
However, GAN-based methods suffer from several challenges. Both AdvGAN and AdvGAN++ generate perturbations in a single iteration, which limits their control over these perturbations. We observe the reason may be the generator continuously increases the perturbations during the iterative process (as illustrated in the Appendix A.1). This may not be effective against adversarial defenses and impacts the attack capability. It is critical since the goal is to maximize attack effectiveness with minimal perturbation. Furthermore, the transferability performance of such attacks is concerning, especially since internal model information is typically unavailable in real-world scenarios.
@@
Inspired by recent works that utilise auto-regressive properties to generate realistic images or text (Chang et al., 2022; Yoon et al., 2019; Ni et al., 2020), we propose a novel GAN-based algorithm, Progressive Auto-Regression AdvGAN (PAR-AdvGAN) to generate adversarial examples with enhanced transferability. PAR-AdvGAN employs a progressive, auto-regressive iterative method to effectively capture the specific structures of input examples. This process gradually generates more diverse and realistic adversarial examples. Specifically, at time step $t$, we combine the input examples $x_{adv}^{t-1}$ from time step $t-1$ with the initial examples $x_{0}$ to generate the perturbation $G(x_{adv}^{t-1},x_{0})$ at time step $t$. Consequently, the manipulated examples shifts from $x+G(x)$ in the original AdvGAN to $x_{adv}^{t-1}+G(x_{adv}^{t-1},x_{0})$ in PAR-AdvGAN.
@@ 2.1 Adversarial Attacks
While numerous adversarial algorithms are dedicated to generating high-quality and robust adversarial samples, gradient-based attack algorithms constitute a main type. FGSM (Goodfellow et al., 2014b) was the first to utilise the model’s gradients, which adds a small perturbation to the input data in the direction of the gradient, thereby maximising the loss function through gradient ascent to achieve optimal attack performance. MI-FGSM (Dong et al., 2018) incorporates a momentum factor in each iteration to mitigate the impact of local optima on the attack success rate. TI-FGSM (Dong et al., 2019) employs shifted images to calculate the input gradient, a process that involves convolving the original image’s input gradient with a kernel matrix.
Other adversarial attack algorithms, such as PGD (Madry et al., 2017), project samples onto suitable attack directions and limit the size of perturbations to generate robust adversarial examples. C&W method minimises the attack’s objective function to optimise the generation process (Carlini & Wagner, 2017). AdvGAN (Xiao et al., 2018) employs an adversarial training process between the generator and discriminator. This process bolsters the generator’s ability to produce adversarial samples, making them challenging for the discriminator to distinguish from genuine data. Besides attack purpose, we also note some other iterative training methods for GANs, such as the Progressive GAN (Karras et al., 2017) which divides the Generator into several layers, with each layer undergoing individual training. In our approach, we consider an auto-regression methods, where each subsequent generation is based on the results of previous step. Although auto-regression GAN has been improved for continuous generation tasks, all we need is the attack result of the last state, so we need to redesign it for this situation.
@@ 2.2 Adversarial Defenses
Adversarial defense represents an effective approach to mitigate the impact of attacks on DNNs. Commonly used adversarial defense techniques include denoising and adversarial training. The denoising technique employs preprocessing mechanisms to filter out adversarial examples, thereby preventing the poisoning of training data and reducing the likelihood of subsequent attacks on the model. Other notable works include HRGD (Liao et al., 2018), R&P (Xie et al., 2017) and so on (Dziugaite et al., 2016; Cohen et al., 2019).
Adversarial training enhances model robustness by incorporating adversarial examples into the training process. Ensemble adversarial training (Hang et al., 2020) works by decoupling the target model from adversarial examples generated by other black-box models, thereby defending against transferable attacks. To enhance the robustness of our algorithm against adversarial defenses, we validated the attack effectiveness of PAR-AdvGAN on the target model subjected to ensemble adversarial training.
@@ 3.2 Perturbation escalation in AdvGAN
Proposition 1 The generator should obtain information about the original sample $x_{0}$.
Proposition 2 To train the generative model, the inputs to the generator should include a significant number of non-initial samples, particularly those encountered during the adversarial process.
Proposition 3 The generator should enforce constraints on the distance between adversarial samples and the initial sample $x_{0}$ throughout the iterative progress.
@@
0:  iteration number $T$, batch size $n$, progressive generator $G$, discriminator $D$, target network $N$, corresponding label $m$, learning rate $\eta_{1}$, $\eta_{2}$, weight hyper-parameter $\lambda_{1},\lambda_{2},\lambda_{3}$
0:  $\theta_{D}$, $\theta_{G}$
@@ 3.3.2 Auto-Regression Iterative Method
To better understand the auto-regression iterative progress, we decompose $\bigtriangledown_{\theta}$ in Eq. 3. Here, $\frac{\partial x+G(x)}{\partial G(x)}$ equals 1, so it can be omitted (See Appendix A.3 for detailed proof). Following this, we further explore the relationship between $\frac{\partial G(x)}{\partial\theta}$ and $\frac{\partial L_{adv}}{\partial x+G(x)}$. Thus, $\frac{\partial G(x)}{\partial\theta}$ represents the degree of change in $G(x)$ with respect to changing in $\theta$. $\frac{\partial L_{adv}}{\partial x+G(x)}$ can be interpreted as the degree of change in $L_{adv}$ when changing $x+G(x)$. Therefore, we can interpret the gradient ascent process of the parameter $\theta$ as a modification of $\theta$ to drive $x+G(x)$ able to obtain a better adversarial effect. At this point, to enable $G$ to iteratively generate perturbations, we will replace $x$ with $x_{adv}^{t}$. This transforms the first part of Eq. 3 into $\frac{\partial L_{adv}}{\partial x_{adv}^{t}+G(x_{adv}^{t})}$, indicating that $G$ continues to generate perturbations based on $x_{adv}^{t}$.
@@
It is worth noting that we did not choose to calculate $L_{D}$ in the form of $log(1-D(x))$ as in AdvGAN. This is because we find that the gradient of $log(1-D(x))$ for $D$ will be very large and not smooth when $D(x)$ is close to 1, and gradient explosion will occur during iteration. We employ a squared form in the Eq. 10 to help mitigate this issue.
@@
It is worth noting that, in untargeted attacks, a larger value of $Cross\ Entropy(x_{adv}^{t},m)$ indicates a more effective adversarial example. Consequently, to enhance the adversarial nature of $x_{adv}^{t}$ during gradient descent on $L_{adv}$, we prepend a negative sign to $Cross\ Entropy(x_{adv}^{t},m)$. It is also feasible to replace $Cross\ Entropy$ with the loss function used in C&W (Carlini & Wagner, 2017).
@@ 4.1.1 Dataset and Models
We conducted the experiments on the ImageNet-compatible dataset consisting of 1000 images with a resolution of 299×299×3 (Papernot et al., 2018)  [^1] https://github.com/cleverhans-lab/cleverhans/tree/master/cleverhans_v3.1.0/examples/nips17_adversarial_competition/dataset. The dataset generation process follows the literature (Dong et al., 2018; 2019; Gao et al., 2021; 2020)
Here we refer to the typical and state-of-the-art transferable adversarial attack methods (Xiao et al., 2018; Xie et al., 2019; Dong et al., 2019; 2018; Lin et al., 2019; Zhang et al., 2022). To ensure experiment fairness, we selected representative models from two types: normally-trained and defense-trained models. The normally trained models include Inceptionv3 (Inc-v3) (Szegedy et al., 2016), Inception-v4 (Inc-v4) (Szegedy et al., 2017), Inception-ResNet-v2 (IncRes-v2) (Szegedy et al., 2017), ResNet-v2-50 (Res-50) (He et al., 2016a; b), ResNet-v2-101 (Res-101) (He et al., 2016a; b), and ResNet-v2-152 (Res-152) (He et al., 2016a; b). As for the defense-trained models through ensemble adversarial training, we selected Inc-v3ens3 (Tramèr et al., 2017), Inc-v3ens4 (Tramèr et al., 2017), and IncResv2ens (Tramèr et al., 2017).
@@ 4.1.2 Baseline Methods
We employ the original AdvGAN (Xiao et al., 2018) algorithm as our baseline to validate the transferability performance by incorporating self-regressive iteration in PAR-AdvGAN. Meanwhile, to evaluate our proposed PAR-AdvGAN, we selected seven state-of-the-art black-box adversarial attack methods as our competitive baselines, including FGSM (Goodfellow et al., 2014b), BIM (Kurakin et al., 2018), PGD (Madry et al., 2017), DI-FGSM (Xie et al., 2019), TI-FGSM (Dong et al., 2019), MI-FGSM (Dong et al., 2018), and SINI-FGSM (Lin et al., 2019).
@@ 4.1.3 Parameter Settings
All experiments in this study are conducted using the Nvidia RTX 6000 Ada 48GB. In all experiments, we set the following fixed parameters for each algorithm according to the settings in Kim (2020). For AdvGAN and PAR-AdvGAN, the training epochs are set to 60. The initial learning rate for both the Generator and Discriminator is set to 0.001, which is then reduced to 0.0001 at the 50th epoch. For DI-FGSM, we set the decay to 0, the resize_rate to 0.9, and the diversity_prob to 0.5. For TI-FGSM, decay is set to 0, kernel_name is set to ”gaussian,” len_kernel is set to 15, resize_rate is set to 0.9, and the diversity_prob is set to 0.5. For MI-FGSM, decay is set to 1. For SINI-FGSM, decay is set to 1, and $m$ is set to 5.
@@ 4.1.4 Metrics
Attack success rate (ASR) is a metric to evaluate the transferability of attacks. It quantifies the average proportion of mislabeled samples among all generated samples after the attack. Thus, a higher attack success rate signifies better transferability. Additionally, we use Frames Per Second (FPS) to assess the attack speed. Another crucial measure, the perturbation rate, is utilised to ensure that the adversarial images do not largely diverge from the original images in visual perception. A low value of this rate suggests that the adversarial examples maintain close visual fidelity to their originals. Detailed formulas are in the Appendix A.4.

### v3

@@ Abstract
Deep neural networks have demonstrated remarkable performance across various domains. However, they are vulnerable to adversarial examples, which can lead to erroneous predictions. Generative Adversarial Networks (GANs) can leverage the generators and discriminators model to quickly produce high-quality adversarial examples. Since both modules train in a competitive and simultaneous manner, GAN-based algorithms like AdvGAN can generate adversarial examples with better transferability compared to traditional methods. However, the generation of perturbations is usually limited to a single iteration, preventing these examples from fully exploiting the potential of the methods. To tackle this issue, we introduce a novel approach named Progressive Auto-Regression AdvGAN (PAR-AdvGAN). It incorporates an auto-regressive iteration mechanism within a progressive generation network to craft adversarial examples with enhanced attack capability. We thoroughly evaluate our PAR-AdvGAN method with a large-scale experiment, demonstrating its superior performance over various state-of-the-art black-box adversarial attacks, as well as the original AdvGAN.Moreover, PAR-AdvGAN significantly accelerates the adversarial example generation, i.e., achieving the speeds of up to 335.5 frames per second on Inception-v3 model, outperforming the gradient-based transferable attack algorithms. Our code is available at: https://github.com/LMBTough/PAR
@@ 1 Introduction
Deep neural networks (DNNs) are widely used in different real-world applications, i.e., image classification [20], emotional analysis [29], and item recommendations [27]. DNNs demonstrate human-surpassed performance when properly trained. However, DNNs can be vulnerable to adversarial examples crafted by attackers [32, 24, 5], which is a concern in safety-critical scenarios. Thus, a practical approach is to develop effective attack algorithms that can assess the robustness of DNNs against adversarial attacks at an early stage, ultimately enhancing model safety.
@@
Currently, both white-box and black-box attack algorithms, such as gradient-based methods like FGSM [10], NAA [38], SSA [23], and optimization-based approaches such as PGD [25] and C&W [1], require continuous computation of the model’s gradient information throughout the attack process. However, they all require extensive running time. Generative Adversarial Networks (GANs) [9] have demonstrated promising results for realistic sample generation by leveraging both generator and discriminator for training [17, 2, 11, 4]. While the generator constructs high-quality examples, a discriminator learn to distinguish the original and generated examples. Furthermore, once the generator is trained, there will be no additional gradient computation for input examples.
@@
As an early GAN-based model, AdvGAN incorporates a perturbation, denoted as $G(x)$, into the original image instance $x$ for attack [34]. AdvGAN aims to obtain the manipulated image $x+G(x)$ from the original instance $x$ through the discriminator. To achieve high attack success rates in both white-box and black-box attacks, AdvGAN introduces an adversarial loss on top of GANs loss, ensuring the adversarial image is generated in a direction more effective for adversarial attacks. Additionally, it employs hinge loss to limit perturbation range, thereby preventing significant deviations between the adversarial and original images. Subsequently, AdvGAN++ [15] further enhances the attack success rate by utilizing latent features instead of input image instances $x$. It optimizes the latent features during adversarial examples generation.
@@
However, GAN-based methods suffer from several challenges. Both AdvGAN and AdvGAN++ generate perturbations in a single iteration, which limits their control over these perturbations. We observe the reason may be the generator continuously increases the perturbations during the iterative process (as illustrated in the Appendix 6.1). This may not be effective against adversarial defenses and impacts the attack capability. It is critical since the goal is to maximize attack effectiveness with minimal perturbation. Furthermore, the transferability performance of such attacks is concerning, especially since internal model information is typically unavailable in real-world scenarios.
@@
Inspired by recent works that utilise auto-regressive properties to generate realistic images or text [2, 37, 26], we propose a novel GAN-based algorithm, Progressive Auto-Regression AdvGAN (PAR-AdvGAN) to generate adversarial examples with enhanced transferability. PAR-AdvGAN employs a progressive, auto-regressive iterative method to effectively capture the specific structures of input examples. This process gradually generates more diverse and realistic adversarial examples. Specifically, at time step $t$, we combine the input examples $x_{adv}^{t-1}$ from time step t1t-1 with the initial examples $x_{0}$ to generate the perturbation $G(x_{adv}^{t-1},x_{0})$ at time step $t$. Consequently, the manipulated examples shifts from $x+G(x)$ in the original AdvGAN to $x_{adv}^{t-1}+G(x_{adv}^{t-1},x_{0})$ in PAR-AdvGAN.
@@ 2.1 Adversarial Attacks
While numerous adversarial algorithms are dedicated to generating high-quality and robust adversarial samples, gradient-based attack algorithms constitute a main type. FGSM [10] was the first to utilise the model’s gradients, which adds a small perturbation to the input data in the direction of the gradient, thereby maximising the loss function through gradient ascent to achieve optimal attack performance. MI-FGSM [6] incorporates a momentum factor in each iteration to mitigate the impact of local optima on the attack success rate. TI-FGSM [7] employs shifted images to calculate the input gradient, a process that involves convolving the original image’s input gradient with a kernel matrix.
Other adversarial attack algorithms, such as PGD [25], project samples onto suitable attack directions and limit the size of perturbations to generate robust adversarial examples. C&W method minimises the attack’s objective function to optimise the generation process [1]. AdvGAN [34] employs an adversarial training process between the generator and discriminator. This process bolsters the generator’s ability to produce adversarial samples, making them challenging for the discriminator to distinguish from genuine data. Besides attack purpose, we also note some other iterative training methods for GANs, such as the Progressive GAN [16] which divides the Generator into several layers, with each layer undergoing individual training. In our approach, we consider an auto-regression methods, where each subsequent generation is based on the results of previous step. Although auto-regression GAN has been improved for continuous generation tasks, all we need is the attack result of the last state, so we need to redesign it for this situation.
@@ 2.2 Adversarial Defenses
Adversarial defense represents an effective approach to mitigate the impact of attacks on DNNs. Commonly used adversarial defense techniques include denoising and adversarial training. The denoising technique employs preprocessing mechanisms to filter out adversarial examples, thereby preventing the poisoning of training data and reducing the likelihood of subsequent attacks on the model. Other notable works include HRGD [21], R&P [35] and so on [8, 3].
Adversarial training enhances model robustness by incorporating adversarial examples into the training process. Ensemble adversarial training [12] works by decoupling the target model from adversarial examples generated by other black-box models, thereby defending against transferable attacks. To enhance the robustness of our algorithm against adversarial defenses, we validated the attack effectiveness of PAR-AdvGAN on the target model subjected to ensemble adversarial training.
@@ 3.2 Perturbation escalation in AdvGAN
Proposition 1. The generator should obtain information about the original sample $x_{0}$.
Proposition 2. To train the generative model, the inputs to the generator should include a significant number of non-initial samples, particularly those encountered during the adversarial process.
Proposition 3. The generator should enforce constraints on the distance between adversarial samples and the initial sample $x_{0}$ throughout the iterative progress.
@@
0: iteration number $T$, batch size $n$, progressive generator $G$, discriminator $D$, target network NN, corresponding label $m$, learning rate $\eta_{1}$, $\eta_{2}$, weight hyper-parameter $\lambda_{1},\lambda_{2},\lambda_{3}$
[…]
@@ 3.3.2 Auto-Regression Iterative Method
To better understand the auto-regression iterative progress, we decompose $\bigtriangledown_{\theta}$ in Eq. 3. Here, $\frac{\partial x+G(x)}{\partial G(x)}$ equals 1, so it can be omitted (See Appendix 6.3 for detailed proof). Following this, we further explore the relationship between $\frac{\partial G(x)}{\partial\theta}$ and $\frac{\partial L_{adv}}{\partial x+G(x)}$. Thus, $\frac{\partial G(x)}{\partial\theta}$ represents the degree of change in $G(x)$ with respect to changing in $\theta$. $\frac{\partial L_{adv}}{\partial x+G(x)}$ can be interpreted as the degree of change in $L_{adv}$ when changing $x+G(x)$. Therefore, we can interpret the gradient ascent process of the parameter $\theta$ as a modification of $\theta$ to drive $x+G(x)$ able to obtain a better adversarial effect. At this point, to enable $G$ to iteratively generate perturbations, we will replace $x$ with $x_{adv}^{t}$. This transforms the first part of Eq. 3 into $\frac{\partial L_{adv}}{\partial x_{adv}^{t}+G(x_{adv}^{t})}$, indicating that $G$ continues to generate perturbations based on $x_{adv}^{t}$.
@@
It is worth noting that we did not choose to calculate $L_{D}$ in the form of log(1D(x))log(1-D(x)) as in AdvGAN. This is because we find that the gradient of log(1D(x))log(1-D(x)) for $D$ will be very large and not smooth when $D(x)$ is close to 1, and gradient explosion will occur during iteration. We employ a squared form in the Eq. 10 to help mitigate this issue.
@@
It is worth noting that, in untargeted attacks, a larger value of $cross\ entropy\ (x_{adv}^{t},m)$ indicates a more effective adversarial example. Consequently, to enhance the adversarial nature of $x_{adv}^{t}$ during gradient descent on $L_{adv}$, we prepend a negative sign to $cross\ entropy(x_{adv}^{t},m)$. It is also feasible to replace crossentropycross\ entropy with the loss function used in C&W [1].
@@ 4.1.1 Dataset and Models
We conducted the experiments on the ImageNet-compatible dataset consisting of 1000 images with a resolution of 299×299×3 [28]  [^1] https://github.com/cleverhans-lab/cleverhans/tree/master/cleverhans_v3.1.0/examples/nips17_adversarial_competition/dataset. The dataset generation process follows the literature [6, 7]
Here we refer to the typical and state-of-the-art transferable adversarial attack methods [34, 36, 7, 6, 22, 38]. To ensure experiment fairness, we selected representative models from two types: normally-trained and defense-trained models. The normally trained models include Inceptionv3 (Inc-v3) [31], Inception-v4 (Inc-v4) [30], Inception-ResNet-v2 (IncRes-v2) [30], ResNet-v2-50 (Res-50) [13, 14], ResNet-v2-101 (Res-101) [13, 14], and ResNet-v2-152 (Res-152) [13, 14]. As for the defense-trained models through ensemble adversarial training, we selected Inc-v3ens3 [33], Inc-v3ens4 [33], and IncResv2ens [33].
@@ 4.1.2 Baseline Methods
We employ the original AdvGAN [34] algorithm as our baseline to validate the transferability performance by incorporating self-regressive iteration in PAR-AdvGAN. Meanwhile, to evaluate our proposed PAR-AdvGAN, we selected seven state-of-the-art black-box adversarial attack methods as our competitive baselines, including FGSM [10], BIM [19], PGD [25], DI-FGSM [36], TI-FGSM [7], MI-FGSM [6], and SINI-FGSM [22].
@@ 4.1.3 Parameter Settings
All experiments in this study are conducted using the Nvidia RTX 6000 Ada 48GB. In all experiments, we set the following fixed parameters for each algorithm according to the settings in [18]. For AdvGAN and PAR-AdvGAN, the training epochs are set to 60. The initial learning rate for both the Generator and Discriminator is set to 0.001, which is then reduced to 0.0001 at the 50th epoch. For DI-FGSM, we set the decay to 0, the resize_rate to 0.9, and the diversity_prob to 0.5. For TI-FGSM, decay is set to 0, kernel_name is set to "gaussian," len_kernel is set to 15, resize_rate is set to 0.9, and the diversity_prob is set to 0.5. For MI-FGSM, decay is set to 1. For SINI-FGSM, decay is set to 1, and $m$ is set to 5.
@@ 4.1.4 Metrics
Attack success rate (ASR) is a metric to evaluate the transferability of attacks. It quantifies the average proportion of mislabeled samples among all generated samples after the attack. Thus, a higher attack success rate signifies better transferability.
Additionally, we use Frames Per Second (FPS) to assess the attack speed. Another crucial measure, the perturbation rate, is utilised to ensure that the adversarial images do not largely diverge from the original images in visual perception. A low value of this rate suggests that the adversarial examples maintain close visual fidelity to their originals. Detailed formulas are in the Appendix 6.4.